Two identical coil springs are mounted side by side so that they jointly support a weight hanger of 2.00 N. When an additional 4.00 N is added to the hanger, it is displaced downward by 5.00 cm. What is the force constant of the individual springs?
Answer #1

force constant of each spring is k.

so when mounted side by side then keq = 2k

and F = keq x

2 = 2k (x1)

x1 = 1/k

for additional weight,

2 + 4 = 2k(x1 + 0.05)

6 = 2kx1 + 2k(0.05)

6 = 2 + 0.01k

k = 40 N/m